OpenStudy (dls):

If f(x)=x^3+ e^(x/2) and g(x)=f inverse x then find g '(1)

OpenStudy (dls):

@experimentX

OpenStudy (experimentx):

use the same technique

OpenStudy (dls):

I did..ill tell you where I am stuck exactly

OpenStudy (experimentx):

put x=1, that's all

OpenStudy (dan815):

swithc y and x and solve for y and u get inverse function

OpenStudy (dls):

\[ \LARGE g'(x)=\frac{1}{f ' g(x))}\] now putting 1 \[ \LARGE g'(1)=\frac{1}{f ' g(1))}\] how do i get the value for f'(g(1))?

OpenStudy (experimentx):

differentiate f(x) and put x = g(x)

OpenStudy (dls):

\[\LARGE f'(x)=3x^2+\frac{1}{2}e^{\frac{x}{2}}\]

OpenStudy (dls):

\[\LARGE f'(g(1))=3(g(1))^2+\frac{1}{2}e^{\frac{g(1)}{2}}\]

OpenStudy (dls):

nowhere near the answer :O

OpenStudy (anonymous):

first off, \(g^{-1}(1)=0\)

OpenStudy (anonymous):

so your job it to compute \[\frac{1}{f'(0)}\]

OpenStudy (dls):

\[\LARGE g'(1)=\frac{1}{3(g(1))^2+\frac{1}{2}e^{\frac{g(1)}{2}}}\] isnt the answer this?

OpenStudy (anonymous):

i.e. take the derivative of \(f\) then compute \(f'(0)\) then take the reciprocal

OpenStudy (anonymous):

hell no

OpenStudy (anonymous):

forget the inverse function, you do not need it

OpenStudy (dls):

@experimentX :O

OpenStudy (anonymous):

these are the steps 1) first notice that \(f^{-1}(1)=0\) 2) compute \(f'\) 3) evaluate \(f'(0)\) 4) take the reciprocal of it

OpenStudy (anonymous):

\[\LARGE g'(1)=\frac{1}{f '( g(1))}=\frac{1}{f '(0)}\]

OpenStudy (dls):

okay f(0)=1 so f inverse 1 =0 f'(x) is already calculated f'(0)=1 i guess :O 1/1=1? :|

OpenStudy (dls):

f'(0)=1/2 sorry

OpenStudy (anonymous):

i think your derivative may be wrong ok now it is right

OpenStudy (anonymous):

so your answer is \(2\)

OpenStudy (anonymous):

$$f(x)=x^3+ e^{x/2}\implies f'(x)=3x^2+\frac12 e^{x/2}\\g(1)=x\implies f(x)=1\\x^3+e^{x/2}=1\implies x=0\\g'(1)=\frac1{f'(0)}=\frac1{3(0)^2+\frac12 e^{0/2}}=\frac1{\frac12}=2$$

OpenStudy (anonymous):

(just supplementary info to go along with satellite73's explanation)

OpenStudy (dls):

thanks everyone :) can anyone tell me what my solution is trying to tell me though?

OpenStudy (experimentx):

g(1) = 0 everything is all right

OpenStudy (dls):

 \[f(\alpha)=1\] \[\alpha^3+e^{\frac{\alpha}{2}}=1\] comparing.. alpha=0 f(0)=1 f inverse 1 =0 i know the final outcome is correct but why did we take alpha n stuff?

OpenStudy (anonymous):

Your solution is the slope of the inverse of \(f\) where \(f=1\).

OpenStudy (anonymous):

@DLS we need to work backwards to figure out where (i.e for what \(x\) or \(\alpha\)) we get \(f=1\). As it turns out, you can solve this one by "inspection" pretty easily.

OpenStudy (dls):

so everytime i substitute x=0 and see the value of f(0) ? what if it gives no clue?

OpenStudy (anonymous):

for one thing, it tells you if you know the slope of \(f\) at a point \((a,b)\), you also know the slope of \(f^{-1}\) at \((b,a)\)

OpenStudy (dls):

how did we get the value for g(1)?

OpenStudy (dls):

\[\Large g(1)=f^{-1}(1)\] so am i supposed to see at what value f gives 1?

OpenStudy (dls):

yuhuuuu @oldrin.bataku

OpenStudy (experimentx):

just plot if you find it hard

OpenStudy (dls):

so we are supposed to do that only?

OpenStudy (experimentx):

yes you needed to find g(1) in your above equation. if inverse can be called function at x=1, then defintely you can find the value of invsere. luckly it turned out to be zero.

OpenStudy (dls):

great! question wont demand anything above 0,1,2 etc until its a very rare case..i hope?

OpenStudy (experimentx):

the question are carefully constructed ... i hope they won't